The linear equation y=0.15x+0.79 represents an estimate of the average cost of gas for year x starting in 1997. The year 1997 would be represented by x = 1, for example, as it is the first year in the study. Similarly, 2005 would be year 9, or x = 9.
a) What year would be represented by x = 4?
b) What x-value represents the year 2018?
c) What is the slope (or rate of change) of this equation?
d) What is the y-intercept?
e) What does the y-intercept represent?
f) Assuming this growth trend continues, what will the price of gasoline be in the year 2018? How did you arrive at your answer?

Answer by mersha(9)

1997, x=1
1998, x=2
1999, x=3
(a) 2000, x=4
(b) 2018, x=22
(c) 0.15x+0.79,rate change is -5.267
(d) y-intercept is 0.79
(e) growth
(f) 0.15(22)+0.79=4.09